OpenStudy (anonymous):

Find the number of decibels for the power of the sound. A rock concert, 5.33 x 10^-6 watts/cm^2 D=? Please show work, so I can understand...ty

OpenStudy (anonymous):

decibels are a measure which relates the intensity of sound at a given level to the minimum sound that can be heard, called the threshold of hearing. That value - the threshold of hearing - is 10^(-12) watts/m^2. The db level of a sound is given by\[\beta=10\log(I/I _{0})\]where Io is the threshold of sound.For your sound intensity the level would be\[\beta=10\log(5.33\times10^{-6}/10^{-12})=10\log(5.33\times10^{6})=10(6.727)=67.3db\]

OpenStudy (anonymous):

sorry should proofread before I click ...relates the intensity of sound at a given level to the INTENSITY OF minimum sound that can be heard